# The Limits of Assumption-free Tests for Algorithm Performance

### v1

@@ 5.4 Algorithm comparison with a stability assumption
In this section, we study the role of stability in algorithm comparison: does the problem CompareAlg become easier if we assume that $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are both stable algorithms? As for algorithm evaluation, here for the problem of algorithm comparison we will draw the same conclusion: aside from extremely high stability (i.e., the consistency regime), adding a stability assumption does not alter the hardness result for testing CompareAlg.
@@
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the comparison function is given by $\psi(y^{\prime},y^{\prime\prime},y)=\ell(y^{\prime},y)-\ell(y^{\prime\prime},y)$, where the loss $\ell$ takes values in $[0,B]$. Let ${\widehat{T}}$ be a black-box test (as in Definition 3), and assume that ${\widehat{T}}$ satisfies the stability-constrained validity condition,
$\begin{split}\mathbb{P}_{P}\left({\widehat{T}}(\mathcal{A}_{0},\mathcal{A}_{1},\mathcal{D}_{N})=1\right)\leq\alpha\textnormal{ for all $\mathcal{A}_{0},\mathcal{A}_{1},P$ such that}\\ \textnormal{ $\beta_{q}(\mathcal{A}_{l},P,n)\leq\gamma_{q}$ for each $l=0,1$ and $\Delta_{n,P}(\mathcal{A}_{0},\mathcal{A}_{1})\leq 0$},\end{split}$ (14)
@@
Then the power of ${\widehat{T}}$ to compare algorithms is bounded as follows: for any $\mathcal{A}_{0},\mathcal{A}_{1}$ and any $P$ with $\beta_{q}(\mathcal{A}_{l},P,n)\leq\gamma_{q}$ for $l=0,1$ and $\Delta_{P,n}(\mathcal{A}_{0},\mathcal{A}_{1})<0$,
@@ 5.4.1 An alternative notion of stability
An alternative way to incorporate stability, which will allow us to work with an arbitrary comparison function $\psi$, is to define stability directly on the pair of algorithms $\mathcal{A}_{0},\mathcal{A}_{1}$. (This type of condition was also considered in (Bayle et al.,, 2020, Section 4), but there they only consider the special case where $\psi$ is still a difference in losses, as in (9).)
@@
${\widehat{f}}_{l,n}=\mathcal{A}_{l}(\{(X_{i},Y_{i})\}_{i\in[n]};\xi),\quad{\widehat{f}}^{-j}_{l,n}=\mathcal{A}_{l}(\{(X_{i},Y_{i})\}_{i\in[n]\backslash\{j\}};\xi),\quad l=0,1.$
@@
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the comparison function $\psi$ takes values in $[-B,B]$. Let ${\widehat{T}}$ be a black-box test (as in Definition 3), and assume that ${\widehat{T}}$ satisfies the stability constrained validity condition,
$\begin{split}\mathbb{P}_{P}\left({\widehat{T}}(\mathcal{A}_{0},\mathcal{A}_{1},\mathcal{D}_{N})=1\right)\leq\alpha\textnormal{ for any $\mathcal{A}_{0},\mathcal{A}_{1},P$ such that}\\ \textnormal{ $\beta_{q}(\mathcal{A}_{0},\mathcal{A}_{1},P,n)\leq\gamma_{q}$ and $\Delta_{n,P}(\mathcal{A}_{0},\mathcal{A}_{1})\leq 0$},\end{split}$ (16)
@@
Then the power of ${\widehat{T}}$ to compare algorithms is bounded as follows: for any $\mathcal{A}_{0},\mathcal{A}_{1}$ and any $P$ with $\beta_{q}(\mathcal{A}_{0},\mathcal{A}_{1},P,n)\leq\gamma_{q}$ and $\Delta_{P,n}(\mathcal{A}_{0},\mathcal{A}_{1})>0$,
@@ 5.5 Relating algorithm comparison and algorithm evaluation
In this paper, we have considered two different problems: comparing the risk of a single algorithm $\mathcal{A}$ against a fixed threshold $\tau$, as in EvaluateAlg, or comparing the risks of two different algorithms $\mathcal{A}_{0},\mathcal{A}_{1}$ against each other, as in CompareAlg. While these problems are clearly related, the hypotheses being tested appear to be distinct, but in this section we will show that the two testing problems are actually the same—in particular, the hypothesis test (11) for algorithm comparison can be viewed as a special case of the test (1) for algorithm evaluation.
@@
as our new loss function. In particular, note that $\tilde{\ell}$ takes values in $[0,2B]$, since $\psi$ takes values in $[-B,B]$.
@@
In other words, the comparison risk $\Delta_{P,n}(\mathcal{A}_{0},\mathcal{A}_{1})$ for the pair $\mathcal{A}_{0},\mathcal{A}_{1}$ is equivalent (up to a transformation) to the evaluation risk $\tilde{R}_{P,n}(\tilde{\mathcal{A}})$. This means that, by replacing the space $\widehat{\mathcal{Y}}$ with $\widehat{\mathcal{Y}}\times\widehat{\mathcal{Y}}$ (and $B$ with $2B$), we can rewrite the algorithm comparison problem as an algorithm evaluation problem: the testing problem (11) for CompareAlg is the same as testing
@@
which is equivalent to the hypothesis test (1) for EvaluateAlg if we take threshold $\tau=B$. Indeed, as we will see in Appendix B.3, the result in Theorem 7 can be easily obtained from the proof of Theorem 4 due to this correspondence.
@@ 6 Discussion
In this paper, we consider the tasks of algorithm evaluation and algorithm comparison. Our main results characterize the limitations of black-box tests to perform inference on these questions, showing that it is impossible for a universally valid test to have high power under any alternative hypothesis, unless the number of available data is far higher than the target sample size. Our results also show that even an assumption of algorithmic stability is not sufficient to circumvent this hardness result, outside of the consistency regime.
Assessing the quality of models and modeling algorithms are common tasks in many fields, we believe that our results are useful in clarifying the gap between questions that assess an algorithm (EvaluateAlg and CompareAlg), versus questions that assess a particular fitted model produced by the algorithm (EvaluateModel and CompareModel). While our results establish the hardness of these first two problems, even under a stability assumption, an important open question for further research is whether there are alternative mild assumptions (on the algorithm and/or on the data) that would enable powerful inference on EvaluateAlg and CompareAlg even under sample size constraints.
@@ Acknowledgements
R.F.B. was supported by the Office of Naval Research via grant N00014-20-1-2337 and by the National Science Foundation via grant DMS-2023109.
Appendix A Proofs: hardness results for algorithm evaluation
A.1 Key lemma: the role of infinite cardinality
@@
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$. Let $\mathcal{A}$ be any algorithm, let $P$ be any distribution on $\mathcal{X}\times\mathcal{Y}$, and let ${\widehat{T}}$ be any black-box test (as in Definition 1). Let random variables $\mathcal{D}_{N},\mathcal{D}^{(1)},\dots,\mathcal{D}^{({\widehat{r}})}$ denote the data sets that appear when sampling data from $P$ and running ${\widehat{T}}$ on algorithm $\mathcal{A}$. Then, for any $\epsilon>0$, there are only finitely many points $(x,y)\in\mathcal{X}\times\mathcal{Y}$ such that [^4] Here we interpret $(x,y)\in\mathcal{D}$, for a data set $\mathcal{D}\in(\mathcal{X}\times\mathcal{Y})^{M}$, to mean that at least one of the $M$ data points in $\mathcal{D}$ is equal to $(x,y)$.
@@
The same result also holds for any algorithms $\mathcal{A}_{0},\mathcal{A}_{1}$, with ${\widehat{T}}$ now denoting a black-box test for comparing the algorithms (as in Definition 3).
@@
First, by definition of a black-box test, ${\widehat{T}}$ runs for a finite number of rounds ${\widehat{r}}$ (note that ${\widehat{r}}$ is a random variable, e.g., it may depend on the outcomes of the previous rounds, and can be arbitrarily large, but must be finite depending on $P$ and $\mathcal{A}$). Find some value $C_{1}$ such that $\mathbb{P}({\widehat{r}}\leq C_{1})\geq 1-\epsilon/4$. Next, for each $r=1,\dots,C_{1}$, the data set $\mathcal{D}^{(r)}$ has a finite size $N_{r}$ (again, $N_{r}$ is a random variable and may be arbitrarily large, but must be finite depending on $P$ and $\mathcal{A}$). For each $r=1,\dots,C_{1}$, find some $C_{2,r}$ such that $\mathbb{P}(N_{r}\leq C_{2,r})\geq 1-\epsilon/4C_{1}$. Let $C_{2}=\max_{r=1,\dots,C_{1}}C_{2,r}$. Therefore, with probability at least $21-\epsilon/$ 2, ${\widehat{T}}$ runs for at most $C_{1}$ many rounds, and each round generates a data set $\mathcal{D}^{(r)}$ with at most $C_{2}$ many data points. Let $\mathcal{E}$ denote this event, so that we have $\mathbb{P}(\mathcal{E})\geq 1-\epsilon/2$. On the event $\mathcal{E}$, then, the combined data set $\mathcal{D}_{N}\cup\mathcal{D}^{(1)}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}$ contains at most $N+C_{1}C_{2}$ many data points. Now consider a data point $(x,y)$ for which
@@
( holds, and $\mathbb{P}(\textnormal{$\mathcal{E}$ holds, and }(x,y)\in\mathcal{D}_{N}\cup\mathcal{D}^{(1)}\cup\dots\cup\mathcal{D}^{({\widehat{r}})})\geq\epsilon-\mathbb{P}(\mathcal{E}^{c})\geq\epsilon/2.$ (19)
@@
This result reveals how the assumption of infinite cardinality will be used—if only finitely many points $(x,y)$ have probability $\geq\epsilon$ of appearing at any point during the black-box test, then we can always find some other data point $(x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}$ whose probability of appearing during the test is $<\epsilon$. Moreover, the proof also suggests how we might derive analogous results for finite $\mathcal{X}\times\mathcal{Y}$ under a computational constraint. Specifically, if ${\widehat{T}}$ is only allowed to call $\mathcal{A}$ a bounded number of times, and if each run of $\mathcal{A}$ is only allowed to input a data set of some bounded size, then these constraints can be used in place of the (arbitrarily large) values $C_{1},C_{2}$ in the proof above.
A.2 Proof of Theorem 1

### v4

@@ 5.4 Algorithm comparison with a stability assumption
In this section, we study the role of stability in algorithm comparison: does the problem CompareAlg become easier if we assume that $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are both stable algorithms? As for algorithm evaluation, here for the problem of algorithm comparison, we will draw the same conclusion: aside from extremely high stability (i.e., the consistency regime), adding a stability assumption does not alter the hardness result for testing CompareAlg.
@@
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the comparison function is given by $\psi(y^{\prime},y^{\prime\prime},y)=\ell(y^{\prime},y)-\ell(y^{\prime\prime},y)$, where the loss $\ell$ takes values in $[0,B]$. Let $^{\widehat{T}}$ be a black-box test (as in Definition 3), and assume that $^{\widehat{T}}$ satisfies the stability-constrained validity condition,
[…]
@@
Then the power of $^{\widehat{T}}$ to compare algorithms is bounded as follows: for any $\mathcal{A}_{0},\mathcal{A}_{1}$ and any $P$ with $\beta_{q}(\mathcal{A}_{l},P,n)\leq\gamma_{q}$ for $l=0,1$ and $\Delta_{P,n}(\mathcal{A}_{0},\mathcal{A}_{1})>0$,
@@ 5.4.1 An alternative notion of stability
An alternative way to incorporate stability, which will allow us to work with an arbitrary comparison function $\psi$, is to define stability directly on the pair of algorithms $\mathcal{A}_{0},\mathcal{A}_{1}$. This type of condition was also considered by Bayle et al., (2020, Section 4).
@@
[…]
@@
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$, and that the comparison function $\psi$ takes values in [B,B][-B,B]. Let $^{\widehat{T}}$ be a black-box test (as in Definition 3), and assume that $^{\widehat{T}}$ satisfies the stability-constrained validity condition,
[…]
@@
Then the power of $^{\widehat{T}}$ to compare algorithms is bounded as follows: for any $\mathcal{A}_{0},\mathcal{A}_{1}$ and any $P$ with $\beta_{q}(\mathcal{A}_{0},\mathcal{A}_{1},P,n)\leq\gamma_{q}$ and $\Delta_{P,n}(\mathcal{A}_{0},\mathcal{A}_{1})>0$,
@@ 5.5 Relating algorithm comparison and algorithm evaluation
In this paper, we have considered two different problems regarding algorithms: comparing the risk of a single algorithm $\mathcal{A}$ against a fixed threshold $\tau$, as in EvaluateAlg, or comparing the risks of two different algorithms $\mathcal{A}_{0},\mathcal{A}_{1}$ against each other, as in CompareAlg. While these problems are clearly related, the hypotheses being tested appear to be distinct, but in this section we will show that the two testing problems are actually the same—in particular, the hypothesis test (14) for algorithm comparison can be viewed as a special case of the test (1) for algorithm evaluation.
@@
as our new loss function. In particular, note that $\tilde{\ell}$ takes values in $[0,2B]$, since $\psi$ takes values in [B,B][-B,B].
@@
[…]
@@
which is equivalent to the hypothesis test (1) for EvaluateAlg if we take threshold $\tau=B$. Indeed, as we will see in Appendix C.3, the result in Theorem 7 can be easily obtained from the proof of Theorem 4 due to this correspondence.
Remark 5.
When the range of the comparison function $\psi$ is unbounded, i.e., $B=\infty$, then similar to the discussion in Section 3.3, we would also expect that assumption-free CompareModel will be impossible. But luckily that in CompareModel or CompareAlg, we have the freedom to choose the comparison function $\psi$. If we choose $\psi$ to be the one in (13), which compares the frequency on which model or algorithm is better, then we have $B=1$. In that case, we would still have that CompareAlg is much harder than CompareModel.
@@ 6 Discussion
In this paper, we consider the tasks of algorithm/model evaluation and algorithm/model comparison. Our main results characterize the difference between these questions and the limitations of black-box tests to perform inference on algorithmic questions, showing that it is impossible for a universally valid test to have high power under any alternative hypothesis, unless the number of available data is far higher than the target sample size. Our results also show that an assumption of algorithmic stability is not sufficient to circumvent this hardness result, unless in the consistency regime.
Assessing the quality of models and modeling algorithms are common tasks in many fields, we believe that our results are useful in clarifying the gap between questions that assess an algorithm (EvaluateAlg and CompareAlg), versus questions that assess a particular fitted model produced by the algorithm (EvaluateModel and CompareModel). While our results establish the hardness of these first two problems, even under a certain stability assumption, an important open question for further research is whether there are alternative mild assumptions (on the algorithm and/or on the data) that would enable powerful inference on EvaluateAlg and CompareAlg even under sample size constraints. There is much room to explore in this direction; here we mention two lines of work, which might provide us with some inspiration. These two lines work aim to achieve distribution-free inference for the algorithm risk in slightly different settings. The first line is conformal risk control. Translate conformal risk control in our context, it assumes we have a class of algorithms parameterized by $\lambda\in\mathbb{R}$, and the loss function is monotone in $\lambda$. Then the main result of conformal risk control (Angelopoulos et al., 2024b, ) shows that we can select one algorithm, i.e., $\hat{\lambda}$, based on our data so that the selected algorithm has risk control without any assumption on the data distribution. Note that compared to our target, conformal risk control relaxes the goal by finding one particular algorithm whose risk is controlled, and it has some requirement on the class of algorithms. The second line of work is about distribution-free excess risk control. Since in many scenarios, the algorithm risk is relative to the problem difficulty, it is natural to compare the target predictor ${\widehat{f}}_{n}$ with the best possible risk achievable via some reference class of functions, say $\mathcal{F}$. For some special $\mathcal{F}$, it has been shown in Mourtada et al., (2022); Mourtada and Gaiffas, (2022) that we can achieve almost distribution-free upper bound for $R_{P,n}(\mathcal{A})-\inf_{f\in\mathcal{F}}R_{P}(f)$—the excess risk of the algorithm $\mathcal{A}$ with respect to $\mathcal{F}$. Compared to our setting, these works choose to relax the target by considering one algorithm, and the inference target is changed to the excess risk.
@@ Acknowledgements
We thank the Editor, the Associated Editor, and two anonymous referees for their helpful suggestions, which helped improve the presentation and quality of this paper. We especially thank one of the reviewers for pointing out the hardness of EvaluateModel when the loss is unbounded, which motivates the discussion in Section 3.3. The authors would also like to thank Lester Mackey for helpful comments on the paper. R.F.B. was supported by the Office of Naval Research via grant N00014-20-1-2337 and by the National Science Foundation via grant DMS-2023109.
Appendix A Inference for Median Risk
Let us define a median version of the same testing problem as in (1) as:
$\texttt{EvaluateAlg}:H_{0}:\,\textnormal{median}(\ell(\mathcal{A}(\mathcal{D}_{n};\xi)(X),Y))\geq\tau\textnormal{ \ versus \ }H_{1}:\,\textnormal{median}(\ell(\mathcal{A}(\mathcal{D}_{n};\xi)(X),Y))<\tau,$ (21)
where $\ell(\mathcal{A}(\mathcal{D}_{n};\xi)(X),Y)$ means that we fit the algorithm $\mathcal{A}$ on $n$ i.i.d. data $\mathcal{D}_{n}$ drawn from $P$ with random seed $\xi$ and then evaluate the loss of the fitted model on the new independent data $(X,Y)\sim P$. Similarly, we can consider the model risk testing problem as
$\texttt{EvaluateModel}:H_{0}:\,\textnormal{median}(\ell({\widehat{f}}_{n}(X),Y))\geq\tau\textnormal{ \ versus \ }H_{1}:\,\textnormal{median}(\ell({\widehat{f}}_{n}(X),Y))<\tau,$ (22)
where ${\widehat{f}}_{n}$ is a fixed model and the only randomness in computing the median is $(X,Y)$. Then, even when the loss function $\ell$ is unbounded, we can verify that solving (22) would be much easier than (21). To see why this is the case, let us denote $\widetilde{\ell}({\widehat{f}}_{n}(X),Y)={\mathds{1}}(\ell({\widehat{f}}_{n}(X),Y)\geq\tau)$, $\widetilde{R}_{P}({\widehat{f}}_{n})=\mathbb{E}_{P}[\widetilde{\ell}({\widehat{f}}_{n}(X),Y)|{\widehat{f}}_{n}]$ and $\widetilde{R}_{P,n}(\mathcal{A})=\mathbb{E}[\widetilde{R}_{P}({\widehat{f}}_{n})]$ where ${\widehat{f}}_{n}=\mathcal{A}(\mathcal{D}_{n};\xi)$. Then we note the testing problems in (21) and (22) are the same as
$\begin{split}\texttt{EvaluateAlg}:H_{0}:\,\widetilde{R}_{P,n}(\mathcal{A})\geq 0.5\textnormal{ \ versus \ }H_{1}:\,\widetilde{R}_{P,n}(\mathcal{A})<0.5,\\ \texttt{EvaluateModel}:H_{0}:\,\widetilde{R}_{P}({\widehat{f}}_{n})\geq 0.5\textnormal{ \ versus \ }H_{1}:\,\widetilde{R}_{P}({\widehat{f}}_{n})<0.5.\end{split}$ (23)
For the new testing problem in (23), the essential loss function is a 0-11 loss, then EvaluateModel is easy as long as NnN-n is large, while the hardness result for EvaluateAlg in Theorem 1 would still apply.
Appendix B Proofs: hardness results for algorithm evaluation
B.1 Key lemma: the role of infinite cardinality
@@
Assume that either $|\mathcal{X}|=\infty$ or $|\mathcal{Y}|=\infty$. Let $\mathcal{A}$ be any algorithm, let $P$ be any distribution on $\mathcal{X}\times\mathcal{Y}$, and let $^{\widehat{T}}$ be any black-box test (as in Definition 1). Let random variables $\mathcal{D}_{N},\mathcal{D}^{(1)},\dots,\mathcal{D}^{({\widehat{r}})}$ denote the data sets that appear when sampling data from $P$ and running $^{\widehat{T}}$ on algorithm $\mathcal{A}$. Then, for any $\epsilon>0$, there are only finitely many points $(x,y)\in\mathcal{X}\times\mathcal{Y}$ such that [^5] Here we interpret $(x,y)\in\mathcal{D}$, for a data set $\mathcal{D}\in(\mathcal{X}\times\mathcal{Y})^{M}$, to mean that at least one of the $M$ data points in $\mathcal{D}$ is equal to $(x,y)$.
@@
The same result also holds for any algorithms $\mathcal{A}_{0},\mathcal{A}_{1}$, with $^{\widehat{T}}$ now denoting a black-box test for comparing the algorithms (as in Definition 3).
@@
First, by definition of a black-box test, $^{\widehat{T}}$ runs for a finite number of rounds $^{\widehat{r}}$ (note that $^{\widehat{r}}$ is a random variable, e.g., it may depend on the outcomes of the previous rounds, and can be arbitrarily large, but must be finite depending on $P$ and $\mathcal{A}$). Find some value $C_{1}$ such that $\mathbb{P}({\widehat{r}}\leq C_{1})\geq 1-\epsilon/4$. Next, for each $r=1,\dots,C_{1}$, the data set $\mathcal{D}^{(r)}$ has a finite size $N_{r}$ (again, $N_{r}$ is a random variable and may be arbitrarily large, but must be finite depending on $P$ and $\mathcal{A}$). For each $r=1,\dots,C_{1}$, find some $C_{2,r}$ such that $\mathbb{P}(N_{r}\leq C_{2,r})\geq 1-\epsilon/4C_{1}$. Let $C_{2}=\max_{r=1,\dots,C_{1}}C_{2,r}$. Therefore, with probability at least $1-\epsilon/2$, $^{\widehat{T}}$ runs for at most $C_{1}$ many rounds, and each round generates a data set $\mathcal{D}^{(r)}$ with at most $C_{2}$ many data points. Let $\mathcal{E}$ denote this event, so that we have $\mathbb{P}(\mathcal{E})\geq 1-\epsilon/2$. On the event $\mathcal{E}$, then, the combined data set $\mathcal{D}_{N}\cup\mathcal{D}^{(1)}\cup\dots\cup\mathcal{D}^{({\widehat{r}})}$ contains at most $N+C_{1}C_{2}$ many data points. Now consider a data point $(x,y)$ for which
@@
$\mathbb{P}(\textnormal{$\mathcal{E}$ holds, and }(x,y)\in\mathcal{D}_{N}\cup\mathcal{D}^{(1)}\cup\dots\cup\mathcal{D}^{({\widehat{r}})})\geq\epsilon-\mathbb{P}(\mathcal{E}^{c})\geq\epsilon/2.$ (25)
@@
This result reveals how the assumption of infinite cardinality will be used—if only finitely many points $(x,y)$ have probability $\geq\epsilon$ of appearing at any point during the black-box test, then we can always find some other data point $(x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}$ whose probability of appearing during the test is $\epsilon$. Moreover, the proof also suggests how we might derive analogous results for finite $\mathcal{X}\times\mathcal{Y}$ under a computational constraint. Specifically, if $^{\widehat{T}}$ is only allowed to call $\mathcal{A}$ a bounded number of times, and if each run of $\mathcal{A}$ is only allowed to input a data set of some bounded size, then these constraints can be used in place of the (arbitrarily large) values $C_{1},C_{2}$ in the proof above.
B.2 Proof of Theorem 1
B.3 Proof of Theorem 3